Grade 12th passMechanics

A boy throws a stone horizontally off a cliff.it hits the sea 2 second later at a distance of 40m from the foot of the cliff calculate the height of the cliff

Adnanullah
5 Years agoGrade 12th pass

1 Answer

REEM HAZAKAT
5 Years ago
Height = 0.5 g t*t= 0.5*10*2*2= 20Initial speed = horizontal dist/ time= 40/ 2 = 20 m/s scalar quantityDirection of stone is in the direction of initial speed if no obstacle is present